# Learning New Tasks from a Few Examples with Soft-Label Prototypes

### v3

@@ Abstract
Existing approaches to few-shot learning in NLP rely on large language models and fine-tuning of these to generalise on out-of-distribution data. In this work, we propose a simple yet powerful approach to “extreme” few-shot learning, wherein models are exposed to as little as $4$ examples per class, based on soft-label prototypes that collectively capture the distribution of different classes across the input domain space. Inspired by previous work Sucholutsky et al. (2021) on univariate or simple multivariate (synthetic) data, we propose a novel approach that is effective on large, high-dimensional and real-world datasets. We learn soft-label prototypes within a neural framework (DeepSLP) and we experimentally demonstrate that it achieves superior performance on 31/48 tested tasks and few-shot settings while closely matching the performance of strong baselines on the rest. We focus on learning previously unseen NLP tasks from very few examples ($4$, $8$, $16$) and present an in-depth analysis of the effectiveness of our approach.
@@ 1 Introduction
Humans have a remarkable ability to adapt knowledge gained in one domain and apply it in another setting, and to identify or disambiguate objects after observing only a handful of examples Lake et al. (2015). This has inspired research in few-shot learning that aims to build models that can also learn a new task using only a small number of examples per class. Early few-shot learning in NLP relied on interventions at the data level, such as dataset augmentation Clark et al. (2018) or generation of adversarial examples from few-shot datasets Miyato et al. (2016). The advent of large language models (LLMs) has now led to a range of approaches, including fine-tuning on different target tasks Sun et al. (2020); Zhou and Srikumar (2022), creating prompt-enhanced few-shot datasets for training Gao et al. (2020); Schick and Schütze (2020); Lester et al. (2021), and the application of meta-learning Finn et al. (2017); Snell et al. (2017) to optimise model parameters such that models adapt quickly to new tasks using past experience Dou et al. (2019); Holla et al. (2020); van der Heijden et al. (2021) which requires a large number of diverse and balanced few-shot meta-training tasks for effective learning.
In this paper, we propose a simple and effective approach to “extreme” few-shot learning where a model generalises on unseen data after adaptation on as little as 4 examples per class, in the absence of any auxiliary training data and without any LLM finetuning. Inspired by Sucholutsky et al. (2021), we first generate compact representations of input data (by solving for linear constraints) called soft-label prototypes that capture the distribution of different classes across the input domain space, and which are used to make predictions. While the original method of Sucholutsky et al. (2021) is designed for small sets of simple univariate or multivariate (synthetic) data, we extend their approach to work on large sets of high-dimensional and real-world NLP datasets. We then propose a novel method that learns such soft-label prototypes within a neural framework, DeepSLP. We focus on learning previously unseen NLP tasks from a few examples per task ($4$, $8$, $16$). We experimentally demonstrate that DeepSLP outperforms a range of strong few-shot baselines in the majority ($31/48$) of investigated tasks, while closely matching their performance on the remaining tasks.
@@ 2 Related work
Few-shot learning approaches in NLP include data augmentation and semi-supervised learning; e.g., augmentation with adversarial examples Miyato et al. (2016), interpolation of training data into a learnable higher dimensional embedding space Chen et al. (2020), and consistency training to make models more resistant to noise Xie et al. (2019). Recent research efforts on large-scale pre-training of language models Devlin et al. (2018); Radford et al. (2019); Brown et al. (2020) reduce the amount of data required for their subsequent fine-tuning in a given task. Instruction tuning Sanh et al. (2021) and in-context learning Liu et al. (2021) show that natural language instructions or prompts can enhance a model’s few-shot learning abilities by leveraging the language (instruction) understanding abilities of the given pretrained LLM. Creating masked language modeling tasks (at training time) with prompts, either in the vocabulary space Gao et al. (2020) or embedding space Li and Liang (2021); Lester et al. (2021), is experimentally demonstrated to give improved performance across classification tasks in few-shot settings compared to LLM fine-tuning for GPT2 Radford et al. (2019) and BART Lewis et al. (2019). Sanh et al. (2021) further report that diverse and additional multi-task prompted training leads to better generalisation, less variation on unseen data and an overall superior performance compared to larger state-of-the-art models. However, the search space over LLMs, prompt templates, and few-shot learning is so great that there is yet to be an established standard.
@@
Previous work has also tackled few-shot learning within the meta-learning paradigm of “learning to learn” , referring to methods that are trained to adapt quickly (in a few gradient steps) and from a small number of examples using past experience Finn et al. (2017). Meta-learning has been applied to a range of tasks, including relation extraction and relation classification Obamuyide and Vlachos (2019a, b); word sense disambiguation Holla et al. (2020); sentiment classification, entity-typing and natural language inference Bansal et al. (2020); Nooralahzadeh et al. (2020); part-of-speech tagging Wang et al. (2020); and dependency parsing Langedijk et al. (2021).
Despite the many successful applications, however, meta-learning comes with complexities with respect to data and computation Antoniou et al. (2019). Such methods typically require a large set of diverse and balanced meta-learning tasks for effective learning Holla et al. (2020), which presents an experimental obstacle to their widespread use, while training times can span several days as their optimisation is computationally expensive Bansal et al. (2020).
Our work is more similar to the few-shot learning approach of Bansal et al. (2019), particularly in terms of tasks and datasets used. However, their approach utilises meta-learning together with auxiliary training data (GLUE; Wang et al. (2018)) and it is therefore not directly comparable to ours that focuses on a more realistic (yet more challenging) few-shot learning setting that does not rely on any additional labelled data. To facilitate a comparison, however, we do present an analysis of model performance when such data is available (Section 9.2).
3 Soft-Label Prototypes (SLP)
@@
Sucholutsky and Schonlau (2021) define a soft label as a vector $Y$ representing a data point’s simultaneous membership to several classes. It is used in cases where we can denote a point’s partial association to different classes. Using this definition, they further define a soft-label prototype as $(\vec{X},Y)$, where $\vec{X}$ is a point in input space (e.g., an input feature vector) and $Y$ is the corresponding soft label. The underlying idea is that a small set of soft-label prototypes can be used to accurately represent a training set.
@@ 3.1 Generating soft-label prototypes
Soft-label prototypes assign soft labels to every point in the input domain; therefore, a soft-label prototype at point $\vec{X}$ represents the class distribution (determined from the training data) at $\vec{X}$. Based on Algorithm 4 in Sucholutsky et al. (2021), the process of generating soft-label prototypes from training data is a two step process — (1) finding lines that connect the class centroids in the data, where each line connects some of the centroids, and every centroid belongs to one line; and (2) deriving soft-label prototypes capturing the class distribution at the ends of each line.
@@ 3.1.1 Finding lines connecting all centroids
The underlying idea here is that we seek to find classes that lie on the same manifold. First, we compute the centroid of each class in the input data. Then, we find and fit class centroids on the minimum number of lines. For this, we use the recursive regression method proposed by Sucholutsky et al. (2021). This method clusters centroids hierarchically to group similar (interval) centroids together, and fits a regression line on the centroids. The similarity of centroids within a single cluster is judged by how well all the centroids fit on a regression line. If the Euclidean distance of a particular centroid is beyond a pre-defined tolerance threshold $\epsilon$ from a line, it is removed from that cluster and assigned to another cluster. We use this method for all our experiments, as we experimentally find (on our development data) that it performs well on high-dimensional data spread across many classes such as the ones we test here. In Figure 0(a), we present an example set of lines connecting all centroids. For further details on recursive regression, we refer the reader to Sucholutsky et al. (2021).
3.1.2 Deriving soft-label prototypes by optimising for linear constraints
Once we find the lines, we use the endpoints of each line as the location of soft-label prototypes. Therefore, for $l$ lines fitted on $n$ centroids we have $2l$ prototypes. Sucholutsky et al. (2021) mathematically define finding the class distribution at each end point / soft-label prototype as a constraint optimisation problem. The optimisation problem consists conceptually of two main sets of constraints: (i) the target class at each centroid has the maximum influence amongst all classes at certain points along the line (endpoint of the line and midpoints between classes); and (ii) the difference between the influence of the target class and the sum of the influences of all other classes along the line is maximised. The output of these optimisation constraints are two soft-label prototypes which are “set” at the ends of each line (i.e., $\vec{X}$) as shown in Figure 0(b). For further details, we refer the reader to Sucholutsky et al. (2021).
3.2 Classification with soft-label prototypes
Given $M$ soft-label prototypes representing the input distribution of $N$ classes, we define $S=(\vec{X_{1}},Y_{1}),...,(\vec{X_{M}},Y_{M})$ to be our set of prototypes, where $\vec{X_{i}}$ is the location of the $i^{th}$ prototype in the input feature space and $Y_{i}$ is a matrix of dimension $[N\times 1]$ denoting the soft labels. Given a test datapoint $\vec{x}$, we find the nearest line to the test point and calculate the Euclidean distances ${D}={(\vec{X_{i}},\vec{x})}_{i=1,2...M}$ from each of the two prototypes of that line to $\vec{x}$. We then sort $S$ in ascending order of distances using ${D}$ and weigh the probability distribution of the $i^{th}$ nearest prototype (with a total of two prototypes) inversely by its distance to the point located at $\vec{x}$ to get $Y^{*}$:
$Y^{*}=\sum_{i=1}^{2}\frac{Y_{i}}{d(\vec{X_{i}},\vec{x})}$ (1)
[…]
3.3 Classification with SLP: A toy example
Figure 2 presents an example classification with SLP. Given the class centroids for blue, green and yellow are located at $(0,0)$, $(1.5,0)$ and $(3,0)$ respectively, two soft-label prototypes are defined by a line connecting yellow and blue, and are thus located at $(3,0)$ and $(0,0)$ respectively. The soft labels in Figure 1(a) contain the per-class probability distribution derived by the SLP method; for example, $p(x=blue)=0.6$ and $p(x=green)=0.4$ for the left prototype, and $p(x=green)=0.4$ and $p(x=yellow)=0.6$ for the right prototype. When a new test instance $x$ located at $(1.5,0.8)$ is presented, we make predictions as follows: we find the nearest line to $x$ and consider its distance from the two prototypes at the ends of the line and multiply the class distribution of each prototype by the inverse distance as per Eq. 1.
Figure 2: Classification example with SLP (figure from Sucholutsky and Schonlau (2021)).
4 Methodology
We encode training and test instances (Section 5) using base cased BERT Devlin et al. (2018) from the HuggingFace Transformers [^1] https://huggingface.co/docs/transformers/index library Wolf et al. (2019). We use the training data to generate soft-label prototypes, and the test data to make predictions as per Eq. 1. We develop and experiment with two variants of the SLP approach, which we describe below. We release our code and models to facilitate further research in the field.
4.1 SLP for NLP data
@@
We use the SLP approach described in Section 3. In order to make SLP powerful enough for large, high-dimensional real-world NLP data (vs univariate, simple multivariate and/or synthetic data), we require an optimiser that scales on such complex data. To this end, we use the MOSEK solver for linear programming MOSEK ApS (2019) in the CVXPY library Diamond and Boyd (2016) to perform the required computations to generate the soft-label prototype class distributions.
4.2 DeepSLP: Learning soft labels via gradient descent
@@
We further develop a novel gradient-based approach to generate soft labels as a function of an input $x$ by minimising training loss on a few-shot dataset. After generating lines connecting all class centroids (Section 3.1.1), we set two soft-label prototypes at the ends of the lines. Each soft-label prototype $p_{i}$ is denoted by $g_{\theta_{i}}(f(x))$ where $g$ is a neural network parameterised by $\theta_{i}$ and $f(x)$ is a point in the input space. The neural network consists of a fixed encoder (BERT) given by $f(x)$, and a trainable linear layer which returns the soft-label probability distribution at any point $x$ given by $g_{\theta_{i}}(x)$. Figure 3 presents a visual representation of our approach. Compared to the “standard” SLP variant (Section 4.1) where we find soft-label probability distributions via linear constraint optimisation, we now parameterise our soft-label probability distribution with a neural network.
Figure 3: Learning soft-label prototypes (soft labels in red) using a trainable linear layer (yellow): example for a 3-class prototype.
The trainable linear layer is trained in a few-shot manner (see data setting in Section 5). Crucially, the encoder parameters are frozen as we need our input data points to have an unchanged location in the input space – changing their position might result in class centroids that were previously lying on a straight line to no longer lie on the line. At prediction time, we utilise both soft-label prototypes along the line using Equation 1. A higher distance between a data point $x$ and a prototype leads to a correspondingly smaller effect of the prototype on the final classification; therefore, we want to penalise the prototype that is closer to $x$ more if there is an incorrect classification. Each prototype is therefore assigned a fraction of the total loss that is proportional to the other prototype’s Euclidean distance from $x$. This way, the closer prototype’s weights are corrected more in case of a misclassification. Our approach (DeepSLP) is outlined in Algorithm 1, while an example optimisation is presented in Figure 4.
Figure 4: Training soft-label prototypes in DeepSLP. Class centroids are represented with large circles that lie on a line (Red, Green, Blue), while training set examples are represented with smaller circles of the same colour. Dotted lines represent the backpropagation error, of which the bolded ones represent a larger error per soft-label prototype (in case of misclassification). Predictions for $x$ are based on the prototypes at each end of the line.
We use cross entropy loss which gives a measure of the difference between the true and predicted labels. We initialise the weights of $f_{\theta}(x)$ using a uniform Xavier initialisation Glorot and Bengio (2010) and use warmup steps to adjust the learning rate. Epochs vary based on the number of classes in the classifier head (between 15 and 25; see datasets used in Section 5): preliminary experiments on the development data show that more epochs are needed when a higher number of classes lie along a line.
@@ 20             $\theta_{ir}\leftarrow\theta_{ir}-\eta\nabla_{\theta_{ir}}loss_{2}$
5 Tasks and datasets

### v4

@@ Abstract
Existing approaches to few-shot learning in NLP rely on large language models (LLMs) and/or fine-tuning of these to generalise on out-of-distribution data. In this work, we propose a novel few-shot learning approach based on soft-label prototypes (SLPs) designed to collectively capture the distribution of different classes across the input domain space. We focus on learning previously unseen NLP tasks from very few examples ($4$, $8$, $16$) per class and experimentally demonstrate that our approach achieves superior performance on the majority of tested tasks in this data-lean setting while being highly parameter efficient. We also show that our few-shot adaptation method can be integrated into more generalised learning settings, primarily meta-learning, to yield superior performance against strong baselines.
Learning New Tasks from a Few Examples with Soft-Label Prototypes
Avyav Kumar Singh1, Ekaterina Shutova2, Helen Yannakoudakis1 1Department of Informatics, King’s College London, United Kingdom 2ILLC, University of Amsterdam, the Netherlands avyav_kumar.singh@kcl.ac.uk, e.shutova@uva.nl, helen.yannakoudakis@kcl.ac.uk
@@ 1 Introduction
Humans have a remarkable ability to adapt knowledge gained in one domain and apply it in another setting, and to identify or disambiguate objects after observing only a handful of examples Lake et al. (2015). This has inspired research in few-shot learning that aims to build models that can learn a new task using only a small number of examples per class. Early few-shot learning in NLP relied on interventions at the data level, such as dataset augmentation Clark et al. (2018) or generation of adversarial examples from few-shot datasets Miyato et al. (2016), while more recent approaches van der Heijden et al. (2021); Langedijk et al. (2022) utilise meta-learning Finn et al. (2017); Snell et al. (2017) to optimise model parameters such that models adapt quickly to new tasks using past experience Dou et al. (2019); Holla et al. (2020); van der Heijden et al. (2021). The advent of large language models (LLMs) has led to a plethora of further methods, including fine-tuning on different target tasks Sun et al. (2020); Zhou and Srikumar (2022), creating prompt-enhanced few-shot datasets for training Gao et al. (2020); Schick and Schütze (2020); Lester et al. (2021) as well as parameter-efficient fine-tuning methods for very large language models, with parameters running into billions Hu et al. (2022); Dettmers et al. (2023).
In this paper, we propose a simple and effective approach to few-shot learning based on soft-label prototypes (SLPs) that capture the distribution of different classes across the input domain space, inspired by previous work on generating compact representations of input training data Sucholutsky et al. (2021). Our contributions are summarised as follows: 1) We develop a novel neural framework for few-shot learning via soft-label prototypes that has a very small computational and memory footprint, and achieves state-of-the-art results in limited-resource settings. Our approach (DeepSLP) does not rely on (expensive) LLM parameter updates or auxiliary training data. 2) We focus on few-shot learning of new, unseen NLP tasks using as little as 4 examples per class, and demonstrate that we outperform strong baselines on the majority of test tasks. 3) We demonstrate that our approach can also be effectively adapted (MetaSLP) in high-resource settings when auxiliary training data is available for few-shot learning, and performs competitively when compared against strong baselines. 4) We release our code and data to facilitate further research in the field. [^1] https://github.com/avyavkumar/meta-learned-lines
@@ 2 Related work
Early few-shot learning approaches in NLP include data augmentation and semi-supervised learning; e.g., augmentation with adversarial examples Miyato et al. (2016), interpolation of training data into a learnable higher dimensional embedding space Chen et al. (2020), and consistency training to make models more resistant to noise Xie et al. (2019). Recent research efforts on large-scale pre-training of language models Devlin et al. (2019); Radford et al. (2019); Brown et al. (2020); Touvron et al. (2023); BigScience Workshop (2023); OpenAI (2024) reduce the amount of data required for their subsequent fine-tuning or utilisation in a given task. Instruction tuning and in-context learning Brown et al. (2020); Gao et al. (2020); Sanh et al. (2021); Liu et al. (2021); Min et al. (2022); Sun et al. (2024); Zhou et al. (2024) show that natural language instructions or prompts can enhance a model’s few-shot learning abilities by leveraging the language (instruction) understanding abilities of the given pretrained LLM Zhao et al. (2021); Liu et al. (2022). To fine-tune extremely large language models (with billions of parameters) efficiently, a host of parameter-efficient fine-tuning techniques have also been developed Hu et al. (2022); Dettmers et al. (2023), which leverage pre-trained LLMs and produce superior results on tasks such as question answering, reasoning, text summarisation, coding, etc. Kotitsas et al. (2024); Jiaramaneepinit et al. (2024); Yang et al. (2024); Ding et al. (2023).
However, the search space over LLMs, prompt templates and few-shot learning is so great that there is yet to be an established standard. Different models require different styles of (few-shot) prompting, and certain prompt templates may work better with specific LLMs and datasets rather than universally across the board (e.g., Davis et al. (2024)). Furthermore, evaluating robustness of state-of-the-art / generative LLMs on new, unseen tasks (OOD generalisation) presents a significant challenge due to their vast and unknown training data, resulting in artificially inflated performance as a result of data leakage Yang et al. (2023).
@@
Previous work has also tackled few-shot learning within the meta-learning paradigm of “learning to learn” Schmidhuber (1987); Bengio et al. (1990); Thrun and Pratt (1998), utilising methods that are trained to adapt quickly (in a few gradient steps) to new tasks and from a small number of examples, using past experience. Meta-learning has emerged as a promising technique for a range of tasks Finn et al. (2017); Koch et al. (2015); Ravi and Larochelle (2017), including NLP such as natural language inference, text classification, etc. Obamuyide and Vlachos (2019a, b); Holla et al. (2020); Bansal et al. (2020b); Nooralahzadeh et al. (2020); Wang et al. (2020); Langedijk et al. (2022); Mueller et al. (2023).
In a similar spirit to parameter-efficient fine-tuning Hu et al. (2022); Dettmers et al. (2023), our work shifts away from the aforementioned paradigms that suffer from lack of standardisation (LLM few-shot prompting) and increased computational complexity for fine-tuning (e.g., fine-tuning extremely large language models such as GPT OpenAI (2024). We present a novel, parameter-efficient few-shot learning framework (DeepSLP) based on soft-label prototypes (SLPs), which we show to be effective on a range of tasks in limited and high-resource settings, while having a substantially smaller computational and memory footprint. While DeepSLP does not rely on LLM fine-tuning or auxiliary training data, we present a variant (MetaSLP) that can be used for few-shot learning via auxiliary data and fine-tuned encoders.
We target few-shot learning of new, unseen tasks (i.e., tasks and classes not previously trained on) (a) in limited-resource settings, without access to auxiliary training data, and with frozen model parameters, and (b) in high-resource settings, with access to auxiliary training data, which are used to update model parameters. For the latter, our work is similar to Bansal et al. (2020a). The authors target few-shot learning of unseen tasks via meta-learning, utilising auxiliary training data.
3 Approach: few-shot learning with Soft-Label Prototypes (SLPs)
@@
A soft label is defined as a vector $Y$ representing a data point’s simultaneous membership to several classes Sucholutsky and Schonlau (2021), essentially denoting a point’s partial association to different classes. Using this definition, a soft-label prototype (SLP) is defined as $(\vec{X},Y)$, where $\vec{X}$ is a point in input space (e.g., an input feature vector) and $Y$ is its corresponding soft label. The underlying idea in Sucholutsky and Schonlau (2021)’s work is that a small set of soft-label prototypes can be used to accurately represent a training set. We build on this idea and reframe SLPs for the task of few-shot learning of new, unseen tasks where very small amounts of data are available per class.
@@ 3.1 Generating soft-label prototypes
Soft-label prototypes assign soft labels to every point in the input domain; therefore, a soft-label prototype at point $\vec{X}$ represents the class distribution (determined from the training data) at $\vec{X}$. The fundamental idea behind a “soft-label” is that, unlike hard labels, which are one-hot encoded labels, soft-labels contain a distribution of probabilistic label values at a particular point in a high-dimensional embedding space.
The process of generating soft-label prototypes from training data can be split into a two step process Sucholutsky et al. (2021): (1) finding lines that connect the class centroids in the training data, where each line connects some of the centroids, and every centroid belongs to one line; and (2) using linear constraints to derive soft-label prototypes capturing the class distribution at the ends of each line. The two steps are presented in detail below.
Figure 1: Learning soft-label prototypes using two trainable linear layers (yellow): example for a 3-class prototype. Dotted lines indicate backpropagation.
@@ 3.1.1 Finding lines connecting all centroids
Here we seek to find classes that lie on the same manifold. First, we compute the centroid of each class in the input training data. Then, we find and fit class centroids on the minimum number of lines using recursive regression Sucholutsky et al. (2021). This method clusters centroids hierarchically to group similar (interval) centroids together, and fits a regression line on the centroids. The similarity of centroids within a single cluster is judged by how well all the centroids fit on a regression line. If the Euclidean distance of a particular centroid is beyond a pre-defined tolerance threshold $\epsilon$ from a line, it is removed from that cluster and assigned to another one. We use this method for all our experiments, as we experimentally find (on our dev data) that it performs well on high-dimensional data spread across many classes such as the ones we test here. In Appendix A.1, Figure 3(a), we present an example set of lines connecting all centroids.
3.1.2 Learning soft-label prototypes
Once we find the lines, we use the endpoints of each line as the location of soft-label prototypes. Therefore, for $l$ lines fitted on $n$ centroids we have $2l$ prototypes. Then, we need to find the class distribution at each end point / soft-label prototype. We develop and experiment with two different approaches to finding the class distributions, one based on constraint optimisation (constraintSLP), and another based on gradient descent (DeepSLP).
Learning via linear constraints (constraintSLP)
@@ 3.1.2 Learning soft-label prototypes
One way in which we can approach this is via constraint optimisation and, specifically, an optimisation problem that consists of two main sets of constraints Sucholutsky et al. (2021): (i) the target class at each centroid has the maximum influence amongst all classes at certain points along the line (endpoint of the line and midpoints between classes); and (ii) the difference between the influence of the target class and the sum of the influences of all other classes along the line is maximised. In order to make this approach powerful enough for large, high-dimensional NLP data, we require an optimiser that scales on such complex data. To this end, we use the MOSEK solver for linear programming MOSEK ApS (2019) in the CVXPY library Diamond and Boyd (2016) to perform the required computations to generate the soft-label prototype class distributions. The output of this is then a set of soft-label prototypes which “sit” at the ends of each line (i.e., $\vec{X}$) as shown in Appendix A.1, Figure 3(b).
Learning via gradient descent (DeepSLP)
@@
Rather than use linear constraints to generate soft-label prototypes, we develop a novel gradient-based approach to generate soft labels as a function of an input $x$ by minimising training loss on a few-shot dataset. After generating lines connecting all class centroids (Section 3.1.1), we set two soft-label prototypes at the ends of the lines. Each soft-label prototype $p_{i}$ is denoted by $g_{i}(f(x))$ where $g$ is a neural network parameterised by $\theta_{i}$ and $f(x)$ is a point in the input space. The neural network consists of a fixed BERT Devlin et al. (2019) encoder [^2] We use BERT as our encoder given its comparatively higher computational efficiency, and do not include LLMs such as Llama and the GPT family as they have already been pre-trained on our test tasks (found here) and hence suffer from data contamination. given by $f(x)$, and a trainable linear layer which returns the soft-label probability distribution at any point $x$ given by $g_{i}(x)$. Figure 1 presents a visual representation of our model. Compared to constraintSLP where we find soft-label probability distributions via linear constraint optimisation, we now parameterise our soft-label probability distribution with a neural network.
@@ 20             $\theta_{ir}\leftarrow\theta_{ir}-\eta\nabla_{\theta_{ir}}loss_{2}$
Crucially, the encoder parameters are frozen as we need our input data points to have an unchanged location in the input space – changing their position might result in class centroids that were previously lying on a straight line to no longer lie on the line. The model is optimised using both soft-label prototypes along a line (see Algorithm 1 below and Section 3.2, Equation 1). Specifically, a higher distance between a data point $x$ and a prototype leads to a correspondingly smaller effect of the prototype on the final classification; therefore, we want to penalise the prototype that is closer to $x$ more if there is an incorrect classification. Each prototype is therefore assigned a fraction of the total loss that is proportional to the other prototype’s Euclidean distance from $x$. This way, the closer prototype’s weights are corrected more in case of a misclassification. The complete algorithm can be seen in Algorithm 1, while an example optimisation is presented in Figure 2.
We use cross entropy loss which gives a measure of the difference between the true and predicted labels. We initialise the weights of $g_{1}(x)$ and $g_{2}(x)$ using a uniform Xavier initialisation Glorot and Bengio (2010) and use warmup steps to adjust the learning rate. Epochs vary based on the number of classes in the classifier head (between 15 and 25; see datasets used in Section 4): preliminary experiments on the development data show that more epochs are needed when a higher number of classes lie along a line.
Figure 2: Training soft-label prototypes in DeepSLP. Class centroids are represented with large circles that lie on a line (Red, Green, Blue), while training set examples are represented with smaller circles of the same colour. Dotted lines represent the backpropagation error, of which the bolded ones represent a larger error per soft-label prototype. Predictions for $x$ are based on the prototypes at each end of the line.
3.2 Classification with soft-label prototypes
Given $M$ soft-label prototypes representing the input distribution of $N$ classes, we define $S=(\vec{X_{1}},Y_{1}),...,(\vec{X_{M}},Y_{M})$ to be our set of prototypes, where $\vec{X_{i}}$ is the location of the $i^{th}$ prototype in the input feature space and $Y_{i}$ is a matrix of dimension $[N\times 1]$ denoting the soft labels. Given a test datapoint $\vec{x}$, we calculate the Euclidean distances ${D}={(\vec{X_{i}},\vec{x})}_{i=1,2...M}$ from each prototype to $\vec{x}$. We then sort $S$ in ascending order of distances using ${D}$, weigh the probability distribution of the $i^{th}$ nearest prototype inversely by its distance to $\vec{x}$, and select the line containing the closest prototype to get $Y^{*}$ Sucholutsky et al. (2021):
$Y^{*}=\sum_{i=1}^{k}\frac{Y_{i}}{d(\vec{X_{i}},\vec{x})}$ (1)
[…]
3.3 Meta-training DeepSLP (MetaSLP)
We further test the suitability of soft-labels in high-resource settings, tuning our text encoder using auxiliary training data. This is similar to the work of Bansal et al. (2020a) that develop a meta-learning approach for few-shot learning of new, unseen tasks while utilising auxiliary training data. We employ a similar approach for rapid generalisation by utilising first-order meta-learning algorithms (which we describe in detail in Appendix A.2). Our model architecture is similar to DeepSLP – it comprises a BERT encoder with two linear layers on top. We train only the last $v$ layers of our encoder to reduce computational overhead, where $v$ is a hyperparameter (See Appendix A.6). We denote the encoder by $f_{\theta}(x)$, and each soft-label prototype at the end of a line by $g_{1}$ and $g_{2}$, parameterised by $\theta_{1}$ and $\theta_{2}$ respectively. The difference between DeepSLP and MetaSLP is that MetaSLP is trained using meta-learning (using auxiliary data described in Section 4), and the encoder $f_{\theta}(x)$ in MetaSLP is fine-tuned (as opposed to being fixed in DeepSLP), following previous work Bansal et al. (2020a).
Inner-loop training
We optimise the soft-label prototypes in the same manner as DeepSLP; i.e., we use Algorithmn 1 to few-shot train the linear layers $g_{1}(x)$ and $g_{2}(x)$ parameterised by $\theta_{1}$ and $\theta_{2}$ respectively. However, meta-learning requires a large set of diverse and balanced meta-learning tasks for effective learning Holla et al. (2020). To ameliorate this, we split the auxiliary datasets (Section 4) used for meta-learning into multiple pairwise tasks to meta-train MetaSLP Bansal et al. (2020a). This means that, during training, we now consider a large number of two-class problems, as opposed to a small number of multi-class problems where the number of classes $n\geq 2$. Such a setting also enables fine-tuning of our encoder $f_{\theta}(x)$. In general, allowing the physical location of encodings to change (in this case via meta-learning’s inner-loop training process), may result in centroids originally connected by a line to no longer be connected by that line (i.e., in the next inner-loop optimisation step). However, if we only meta-train on tasks that focus on two classes at a time, this can trivially ensure that the same line is utilised each time. Our inner-loop optimisation process is given in Algorithm 3 in Appendix A.4.
Outer-loop training
We perform meta-learning using the updated parameters in the inner-loop training process. We experiment with both Reptile Nichol et al. (2018) and FOMAML Finn et al. (2017) as our meta-learning algorithms. Reptile can be considered a simpler variant to MAML-based meta-learning algorithms. We present our outer-loop process in Algorithm 4, Appendix A.4.
Meta-testing
We construct lines for the test sets using the trained MetaSLP model, and fine-tune them on the few-shot adaptation training data for each test task. We then use these lines for classification as described in Section 3.2.
4 Experimental settings and datasets
Experimental settings
We experiment with two settings in terms of amounts of available data. The first is a limited-resource setting where we only train / fine-tune our models in a few-shot manner on small amounts of training data (i.e., in the absence of auxiliary training data). The other setting is a high-resource setting where we assume that auxiliary training data is available for additional training / fine-tuning.
Datasets
We tackle few-shot learning of previously unseen tasks (i.e., not seen during training/fine-tuning), and so our work is similar to Bansal et al. (2020a). For our high-resource setting, we train and test our models on the same data as Bansal et al. (2020a) to ensure direct comparability. For our limited-resource setting, we test in the same way but do not utilise any auxiliary training data; i.e., we only utilise few-shot fine-tuning data for unseen tasks (i.e., only using a very small set of training/fine-tuning examples for the test tasks).
High-resource setting auxiliary training data
Similar to Bansal et al. (2020a), we use GLUE (Wang et al., 2018) to train our models in the high-resource setting. This dataset consists of a range of natural language tasks such as entailment, classification and textual similarity, which are used for model training and evaluation. We use only the training split for meta-learning. Similar to Bansal et al. (2020a), the MNLI Williams et al. (2018) and SNLI Bowman et al. (2015) entailment tasks, which are three-label classification problems, are split in a pairwise manner such that they are included as multiple two-label datasets during training. Following Bansal et al. (2020a), we also train for detecting the sentiment contained within phrases of a sentence by using the phrase-level annotations in SST2 Wang et al. (2018). We utilise the same validation sets – labelled Amazon review data from music, toys and videos for sentiment classification Blitzer et al. (2007). We provide dataset and training details in Appendix A.5.
Evaluation data